Solve the following differential equation:

x^2 d^2y/dx^2 - x dy/dx + 4y = cos(logx) + x sin(logx)
1
Expert's answer
2021-05-11T09:01:10-0400

"x^2 y''-xy'+4y=cos(logx)+xsin(logx)\\
x=e^t, y'=e^{-t}y'_t , y''=e^{-2t}(y''_t-y'_t)\\
e^{2t}e^{-2t}(y''_t-y'_t)-e^te^{-t}y'_t +4y=cost+e^tsint\\
y''_t-2y'_t+4y=cost+e^tsint"

"1) y''_t-2y'_t+4y=0\\
\lambda^2 -2\lambda+4=0\\
D=4-16=-12=12i^2\\
\lambda_1=\frac{2-2\sqrt3 i}{2}=1-\sqrt3 i\\
\lambda_2=\frac{2+2\sqrt3 i}{2}=1+\sqrt3 i\\
y_1=C_1e^tcos\sqrt3 t+C_2e^tsin\sqrt3 t"

"2) a) f_1(t)=cost\\
y_2=acost+bsint\\
y'_2=-asint+bcost\\
y''_2=-acost-bsint\\
-acost-bsint+2asint-2bcost+4acost+4bsint=cost\\
cost:-a-2b+4a=1\\
sint:-b+2a+4b=0\\
3a-2b=1\\
2a+3b=0\\
a=\frac{3}{13}, b=-\frac{2}{13}\\
y_2=\frac{3}{13}cost-\frac{2}{13}sint"

"b) f_2(t)=e^tsint\\
y_3=ae^tcost+be^tsint=e^t(acost+bsint)\\
y'_3=e^t(acost+bsint)+e^t(-asint+bcost)=\\
=e^t((a+b)cost+(b-a)sint)\\
y''_3=e^t((a+b)cost+(b-a)sint)+\\
+e^t(-(a+b)sint+(b-a)cost)=\\
=e^t(2bcost-2asint)"

"e^t(2bcost-2asint)-2e^t((a+b)cost+(b-a)sint)+\\
+4e^t(acost+bsint)=e^tsint\\
cost:2b-2(a+b)+4a=0\\
sint:-2a-2(b-a)+4b=1\\
a=0\\ b=\frac{1}{2}\\
y_3=\frac{1}{2}e^tsint\\"

"y=y_1+y_2+y_3=\\
=C_1e^tcos\sqrt3 t+C_2e^tsin\sqrt3 t+\\
+\frac{3}{13}cost-\frac{2}{13}sint+\frac{1}{2}e^tsint"

Return to "x"

"y=C_1xcos(\sqrt3 log|x|)+C_2xsin(\sqrt3 log|x|)+\\
+\frac{3}{13}cos(log|x|)-\frac{2}{13}sin(log|x|)+\frac{1}{2}xsin(log|x|)"